Given:  C = 100 + 0.75Yd;  I = 120 - 0.1Y – 200i; G= 40; T = 40 

What is the autonomous spending multiplier?


1
Expert's answer
2021-06-15T12:44:54-0400

The multiplier reflects how the change in expenditure would affect total expenditure and the aggregate demand in the country.

Since the change in the consumption expenditure determines the multiplier. Hence marginal propensity to consume would determine the multiplier. 

Marginal propensity to consume refers to the slope of the consumption function.

Marginal propensity to consume refers to the change in the consumption expenditure when there is a change in the disposable income. 


"Multiplier=\frac{1}{1-Marginal\space propensity\space to\space consume}"


"=\frac{1}{1-0.75}"


"=\frac{1}{0.25}"


=4


The multiplier is equal to 4. 

The multiplier can also be calculated by the ratio of one and marginal propensity to save. 

Since marginal propensity to save is equal to 0.25 (i.e., =1-MPC), therefore the multiplier is equal to 4.


The autonomous spending multiplier is equal to 4.